# A Class of Generalized Shift-Splitting Preconditioners for Double Saddle Point Problems

### v3

@@ 5 Two relaxed variants of GSS preconditioner
Assume that $A$ and $D$ are SPD matrices, $B$ and $C$ have full row rank, and let $Q$ and $R$ be SPD matrices. Then, the preconditioned matrix $\mathcal{M}_{\mathrm{RGSS}{\text{-}}\mathrm{I}}^{-1}\mathcal{B}$ has ${\displaystyle{\frac{1}{\bm{\omega}}}}$ as the eigenvalue with multiplicity $n.$ Further, let $\mu$ be an eigenvalue among the remaining $m+l$ eigenvalues with the corresponding eigenvector $[u^{T},v^{T}]^{T}$ such that $\|\sqrt{\beta}Q^{1/2}u^{T},\sqrt{\tau}R^{1/2}v^{T}]^{T}\|_{2}=1$. Then
@@
Let $S=BA^{-1}B^{T}.$ Then the saddle point matrix $\mathcal{B}$ and the preconditioner $\mathcal{M}_{\text{RGSS-II}}$ admits the following decompositions:
$\mathcal{B}=\mathbf{L}\mathbf{D}\mathbf{U}\quad\text{and}\quad\mathcal{M}_{\text{RGSS-I}}=\mathbf{L}\widetilde{\mathbf{D}}\mathbf{U},$ (5.11)
where $\mathbf{L}=\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&\bm{0}&I\end{matrix}\right],$ $\mathbf{D}=\left[\begin{matrix}A&\bm{0}&\bm{0}\\ \bm{0}&D&C\\ \bm{0}&-C^{T}&S\end{matrix}\right],$ $\widetilde{\mathbf{D}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\bm{\omega}D+\beta Q&\bm{\omega}C\\ \bm{0}&-\bm{\omega}C^{T}&S+\tau R\end{matrix}\right]$ and $\mathbf{U}=\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&\bm{0}\\ \bm{0}&\bm{0}&I\end{matrix}\right].$ Using decompositions in (5.11), we obtain
$\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}=\mathbf{U}^{-1}\widetilde{\mathbf{D}}^{-1}\mathbf{D}\mathbf{U}.$ (5.12)
Therefore, $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ is similar to $\widetilde{\mathbf{D}}^{-1}\mathbf{D},$ which is given by
@@
where $\mathbf{M}=\left[\begin{matrix}\bm{\omega}D+\beta Q&\bm{\omega}C\\ -\bm{\omega}C^{T}&\bm{\omega}S+\tau R\end{matrix}\right]$ and $\mathbf{A}=\left[\begin{matrix}D&C\\ -C^{T}&S\end{matrix}\right].$ Hence, $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ has an eigenvalue ${\displaystyle{\frac{1}{\bm{\omega}}}}$ with multiplicity at least $n$ and while the remaining eigenvalues are those of the preconditioned matrix $\mathbf{M}^{-1}\mathbf{A}.$ Consider $\mathbf{I}=\left[\begin{matrix}\bm{0}&I\\ -I&\bm{0}\end{matrix}\right]\in{\mathbb{R}}^{l+m},$ then $\mathbf{I}\left(\mathbf{I}^{-1}\mathbf{M}\mathbf{I}\right)^{-1}\left(\mathbf{I}^{-1}\mathbf{A}\mathbf{I}\right)\mathbf{I}^{-1}.$ Consequently, $\mathbf{M}^{-1}\mathbf{A}$ is similar to $\widetilde{\mathbf{M}}^{-1}\widetilde{\mathbf{A}},$ where
@@
The following theorem establishes the spectral properties for the RGSS-II preconditioned matrix $\mathcal{M}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$.
@@
Assume that $A$ and $D$ are SPD matrices, $B$ and $C$ have full row rank, and let $Q$ be an SPD matrix. Then the preconditioned matrix $\mathcal{M}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$ has the eigenvalue ${\displaystyle{\frac{1}{\bm{\omega}}}}$ with multiplicity $n+l.$ The remaining eigenvalues satisfies the generalized eigenvalue problem $(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R){\bf x}=\lambda\mathcal{R}_{2}{\bf x},$ where $\mathcal{R}_{2}=\tau R+\bm{\omega}BA^{-1}B^{T}+\bm{\omega}C^{T}D^{-1}C.$
@@
$\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ $=\left[\begin{matrix}\bm{\omega}^{-1}A&\bm{0}&-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\\ \bm{0}&\bm{\omega}^{-1}D&-D^{-1}C\mathcal{R}_{2}^{-1}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}\end{matrix}\right]\left[\begin{matrix}A&\bm{0}&B^{T}\\ \bm{0}&D&C\\ \bm{0}&\bm{0}&\mathbf{X}\end{matrix}\right]$
@@
where $\mathbf{X}=BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R.$ Thus, $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ has the eigenvalue ${\displaystyle{\lambda=\frac{1}{w}}}$ with multiplicity at least $n+l,$ and the rest of eigenvalues satisfy the generalized eigenvalue problem
@@
Suppose that assumption on Theorem 5.2 holds. Then the eigenvalues of $\mathcal{M}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$ satisfies
@@
Next, we will examine the properties of the minimal polynomial of the preconditioned matrix $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B},$ which is crucial for effective Krylov subspace acceleration.
@@
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row rank matrices. Then, the degree of the minimal polynomial of the preconditioned matrix $\mathcal{M}_{\mathrm{RGSS-II}}^{-1}\mathcal{B}$ is almost $m+1.$ Therefore, the dimension of the Krylov subspace $\mathcal{K}(\mathcal{M}_{\mathrm{RGSS-II}}^{-1}\mathcal{B},{\bm{b}})$ is almost $m+1.$
@@
$\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}=\left[\begin{matrix}w^{-1}I&\bm{0}&\Sigma_{1}\\ \bm{0}&w^{-1}I&\Sigma_{2}\\ \bm{0}&\bm{0}&\Sigma_{3}\end{matrix}\right],$ (5.20)
where $\Sigma_{1}=\bm{\omega}^{-1}AB^{T}-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\mathbf{X},$ $\Sigma_{2}=\bm{\omega}^{-1}DC-D^{-1}C\mathcal{R}_{2}^{-1}\mathbf{X},$ $\Sigma_{3}=\mathcal{R}_{2}^{-1}\mathbf{X},$ and $\mathbf{X}=BA^{-1}B^{T}+C^{T}D^{-1}C+R.$ Let $\mu_{i},$ $i=1,2,\ldots,m$ be the eigenvalues of of $\Sigma_{3}.$ Then they are also eigenvalues of the preconditioned matrix $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}.$ Then the characteristic polynomial of $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ is given by
$f(\lambda)=\left(\lambda-\frac{1}{w}\right)^{n+l}\prod_{i=1}^{m}(\lambda-\mu_{i}).$
@@
$g(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B})$ $=(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B}-\frac{1}{\bm{\omega}}I)\prod\limits_{i=1}^{m}(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B}-\mu_{i}I)$
@@
Given that $\Sigma_{3}$ has the eigenvalues $\mu_{i},$ $i=1,2,\ldots,m,$ we obtain ${\prod\limits_{i=1}^{m}(\Sigma_{3}-\mu_{i}I)}=\bm{0}.$ Therefore, from (5.21), we obtain $g(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B})=\bm{0}.$ Hence, by Cayley-Hamilton theorem, the degree of the minimal polynomial of $\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B}$ is almost $m+1.$
As mentioned in [35], the degree of the minimal polynomial of a matrix and the dimension of the associated Krylov subspace are equal. Therefore, the Krylov subspace $\mathcal{K}(\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B},{\bm{b}})$ has dimension atmost $m+1.$ Hence, the proof follows. $\blacksquare$
@@ 6 Discussion on the selection of the parameters
It is worth noting that proposed GSS, RGSS-I and RGSS-II preconditioners involve the parameters $\alpha,$ $\beta,$ $\tau$ and $\bm{\omega}.$ As a general criterion for a preconditioner to perform efficiently, it should be as close as possible to the coefficient matrix of the system [10], we find the optimal choices for the proposed GSS preconditioner by minimizing $\|\mathcal{N}_{\text{GSS}}\|_{F}=\|\mathcal{M}_{\text{GSS}}-\mathcal{B}\|_{F}.$ First, we define the function $\varphi$ by
@@
In the sequel, the GSS preconditioner is equivalently rewritten as: $\mathcal{M}_{\text{GSS}}=\bm{\omega}\widetilde{\mathcal{M}}_{\text{GSS}},$ where
~$\widetilde{\mathcal{M}}_{\text{GSS}}=\left[\begin{matrix}\frac{\alpha}{\bm{\omega}}P+A&\bm{0}&B^{T}\\ \bm{0}&\frac{\beta}{\bm{\omega}}Q+D&C\\ -B&-C^{T}&\frac{\tau}{\bm{\omega}}R\end{matrix}\right]$
can be regarded as a scaled preconditioner. A preconditioner is considered efficient if it closely approximates the coefficient matrix, and for given $\alpha,\beta,\tau>0,$$\widetilde{\mathcal{M}}_{\text{GSS}}-\mathcal{B}=\frac{1}{\bm{\omega}}\Theta\rightarrow\bm{0}$ as $\bm{\omega}\rightarrow\infty.$ Similar studies apply to the RGSS-I and RGSS-II preconditioners as well. The performance of the proposed preconditioners by varying the parameters is shown in the numerical experiment section.
@@ 7 Numerical experiments
To demonstrate the effectiveness and robustness of the proposed preconditioners GSS, RGSS-I and RGSS-II over the existing ones within the Krylov subspace iterative method to solve the DSPP, in this section, we perform a few numerical experiments. We compare our proposed GSS, RGSS-I and RGSS-II PGMRES methods (abbreviated as “GSS”, “RGSS-I” and “‘RGSS-II”, respectively) with the GMRES method and PGMRES methods with BD [13], DS [8], RDF [11], and SS [18] preconditioners (abbreviated as “BD”, “DS”, “RDF”, “SDF” and “SS”, respectively). The numerical results are presented in terms of iteration counts (abbreviated as “IT”) and elapsed CPU time in seconds (abbreviated as “CPU”).
@@
or if the maximum number of iterations exceeds $5000$. All the linear subsystems involved in Algorithms 1, 2 and 3 are solved using the Cholesky Factorization. Numerical experiments are conducted in MATLAB R2024a on a Windows 11 system, using an Intel(R) Core(TM) i7-10700 CPU at 2.90 GHz with 16 GB of memory.
@@
$\begin{array}[]{cc}&\min_{u,f}\frac{1}{2}\|u-\hat{u}\|^{2}_{L_{2}(\Omega)}+\frac{\bm{\beta}}{2}\|f\|^{2}_{L_{2}(\Omega)}\\ \text{such that }&-\nabla^{2}u=f\leavevmode\nobreak\ \text{in}\leavevmode\nobreak\ \Omega,\\ &\quad\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u=g\leavevmode\nobreak\ \text{on}\leavevmode\nobreak\ \partial\Omega,\end{array}$ (7.4)
where $u$ is the state, and $\hat{u}$ is the desired state, $f$ is the control, $\Omega=[0,1]\times[0,1]$ is the domain with the boundary $\partial\Omega,$ and $0<{\bm{\beta}}\ll 1$ is the regularization parameter. By discretizing (7.4) using the Galerkin finite element method and then applying Lagrange multiplier techniques, we obtain the following system:
$\left[\begin{matrix}{\bm{\beta}}M&\bm{0}&K^{T}\\ \bm{0}&M&-M\\ -K&M&\bm{0}\end{matrix}\right]\left[\begin{matrix}u\\ f\\ \bm{\lambda}\end{matrix}\right]=\left[\begin{matrix}\bm{0}\\ b\\ -d\end{matrix}\right],$ (7.5)
where $M\in{\mathbb{R}}^{n\times n}$ and $K\in{\mathbb{R}}^{n\times n}$ are SPD mass matrix and discrete Laplacian, respectively. Note that by setting $A=\bm{\beta}M$, $B=K$, $C=-M$, and $\mathbf{b}=[\mathbf{0},b^{T},-d^{T}]^{T}$, (7.5) can be expressed in the form of the DSPP (1.1), where $n=m=p$.
The MATLAB code downloaded from [31], generates the linear system of the form (7.5) by using the following setup: parameters in “set_def_setup.m” are selected as def_setup.bc=‘dirichlet’, def_setup.beta=$1e-2$, def_setup.ob=1, def_setup.type=‘dist2d’ and def_setup.pow = $5$, $6$ and $7$.
Table 1: Experimental results of GMRES, BD, DS, RDF, SS, GSS, RGSS-I and RGSS-II PGMRES methods for Example 1 when ${\bm{\beta}}=0.1$
@@
Table 2: Experimental results of GMRES, BD, DS, RDF, SS, GSS, RGSS-I and RGSS-II PGMRES methods for Example 1 when ${\bm{\beta}}=0.001$

### v4

@@ 5 Two relaxed variants of GSS preconditioner
Assume that $A$ and $D$ are SPD matrices, $B$ and $C$ have full row rank, and let $Q$ and $R$ be SPD matrices. Then, the preconditioned matrix $\mathcal{P}_{\mathrm{RGSS}{\text{-}}\mathrm{I}}^{-1}\mathcal{B}$ has ${\displaystyle{\frac{1}{\bm{\omega}}}}$ as the eigenvalue with multiplicity $n.$ Further, let $\mu$ be an eigenvalue among the remaining $m+l$ eigenvalues with the corresponding eigenvector $[u^{T},v^{T}]^{T}$ such that $\|\sqrt{\beta}Q^{1/2}u^{T},\sqrt{\tau}R^{1/2}v^{T}]^{T}\|_{2}=1$. Then
@@
Let $S=BA^{-1}B^{T}.$ Then the saddle point matrix $\mathcal{B}$ and the preconditioner $\mathcal{P}_{\text{RGSS-I}}$ admits the following decompositions:
$\mathcal{B}=\mathbf{L}\mathbf{D}\mathbf{U}\quad\text{and}\quad\mathcal{P}_{\text{RGSS-I}}=\mathbf{L}\widetilde{\mathbf{D}}\mathbf{U},$ (5.11)
where
$\mathbf{L}$ $=\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&\bm{0}&I\end{matrix}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{D}=\left[\begin{matrix}A&\bm{0}&\bm{0}\\ \bm{0}&D&C\\ \bm{0}&-C^{T}&S\end{matrix}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \widetilde{\mathbf{D}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\bm{\omega}D+\beta Q&\bm{\omega}C\\ \bm{0}&-\bm{\omega}C^{T}&S+\tau R\end{matrix}\right]$
$\text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{U}=\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&\bm{0}\\ \bm{0}&\bm{0}&I\end{matrix}\right].$
Using decompositions in (5.11), we obtain
$\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}=\mathbf{U}^{-1}\widetilde{\mathbf{D}}^{-1}\mathbf{D}\mathbf{U}.$ (5.12)
Therefore, $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ is similar to $\widetilde{\mathbf{D}}^{-1}\mathbf{D},$ which is given by
@@
where $\mathbf{M}=\left[\begin{matrix}\bm{\omega}D+\beta Q&\bm{\omega}C\\ -\bm{\omega}C^{T}&\bm{\omega}S+\tau R\end{matrix}\right]$ and $\mathbf{A}=\left[\begin{matrix}D&C\\ -C^{T}&S\end{matrix}\right].$ Hence, $\mathcal{P}_{\text{RGSS-I}}^{-1}\mathcal{B}$ has an eigenvalue ${\displaystyle{\frac{1}{\bm{\omega}}}}$ with multiplicity at least $n$ and while the remaining eigenvalues are those of the preconditioned matrix $\mathbf{M}^{-1}\mathbf{A}.$ Consider $\mathbf{I}=\left[\begin{matrix}\bm{0}&I\\ -I&\bm{0}\end{matrix}\right]\in{\mathbb{R}}^{l+m},$ then $\mathbf{M}^{-1}\mathbf{A}=\mathbf{I}\left(\mathbf{I}^{-1}\mathbf{M}\mathbf{I}\right)^{-1}\left(\mathbf{I}^{-1}\mathbf{A}\mathbf{I}\right)\mathbf{I}^{-1}.$ Consequently, $\mathbf{M}^{-1}\mathbf{A}$ is similar to $\widetilde{\mathbf{M}}^{-1}\widetilde{\mathbf{A}},$ where
@@
The following theorem establishes the spectral properties for the RGSS-II preconditioned matrix $\mathcal{P}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$.
@@
Assume that $A$ and $D$ are SPD matrices, $B$ and $C$ have full row rank, and let $R$ be an SPD matrix. Then the preconditioned matrix $\mathcal{P}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$ has the eigenvalue ${\displaystyle{\frac{1}{\bm{\omega}}}}$ with multiplicity $n+l.$ The remaining eigenvalues satisfy the generalized eigenvalue problem $(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R){\bf x}=\lambda\mathcal{R}_{2}{\bf x},$ where $\mathcal{R}_{2}=\tau R+\bm{\omega}BA^{-1}B^{T}+\bm{\omega}C^{T}D^{-1}C.$
@@
$\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ $=\left[\begin{matrix}\bm{\omega}^{-1}A&\bm{0}&-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\\ \bm{0}&\bm{\omega}^{-1}D&-D^{-1}C\mathcal{R}_{2}^{-1}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}\end{matrix}\right]\left[\begin{matrix}A&\bm{0}&B^{T}\\ \bm{0}&D&C\\ \bm{0}&\bm{0}&\mathbf{X}\end{matrix}\right]$
@@
where $\mathbf{X}=BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R.$ Thus, $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ has the eigenvalue ${\displaystyle{\lambda=\frac{1}{\bm{\omega}}}}$ with multiplicity at least $n+l,$ and the rest of eigenvalues satisfy the generalized eigenvalue problem
@@
Suppose that assumption on Theorem 5.2 holds. Then, the eigenvalues of $\mathcal{P}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$ satisfy
@@
Next, we will examine the properties of the minimal polynomial of the preconditioned matrix $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B},$ which determines the dimension of the Krylov subspace.
@@
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are nonsingular matrices with $A_{H}$ and $D_{H}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row rank matrices. Then, the degree of the minimal polynomial of the preconditioned matrix $\mathcal{P}_{\mathrm{RGSS-II}}^{-1}\mathcal{B}$ is at most $m+1.$ Therefore, the dimension of the Krylov subspace $\mathcal{K}(\mathcal{P}_{\mathrm{RGSS-II}}^{-1}\mathcal{B},{\bm{b}})$ is at most $m+1.$
@@
$\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}=\left[\begin{matrix}\bm{\omega}^{-1}I&\bm{0}&\Sigma_{1}\\ \bm{0}&\bm{\omega}^{-1}I&\Sigma_{2}\\ \bm{0}&\bm{0}&\Sigma_{3}\end{matrix}\right],$ (5.20)
where $\Sigma_{1}=\bm{\omega}^{-1}AB^{T}-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\mathbf{X},$ $\Sigma_{2}=\bm{\omega}^{-1}DC-D^{-1}C\mathcal{R}_{2}^{-1}\mathbf{X},$ $\Sigma_{3}=\mathcal{R}_{2}^{-1}\mathbf{X},$ and $\mathbf{X}=BA^{-1}B^{T}+C^{T}D^{-1}C+R.$ Let $\mu_{i},$ $i=1,2,\ldots,m$ be the eigenvalues of of $\Sigma_{3}.$ Then they are also eigenvalues of the preconditioned matrix $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}.$ Then the characteristic polynomial of $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ is given by
$f(\lambda)=\left(\lambda-\frac{1}{\bm{\omega}}\right)^{n+l}\prod_{i=1}^{m}(\lambda-\mu_{i}).$
@@
$g(\mathcal{P}_{\text{GSS-II}}^{-1}\mathcal{B})$ $=(\mathcal{P}_{\text{GSS-II}}^{-1}\mathcal{B}-\frac{1}{\bm{\omega}}I)\prod\limits_{i=1}^{m}(\mathcal{P}_{\text{GSS-II}}^{-1}\mathcal{B}-\mu_{i}I)$
@@
Given that $\Sigma_{3}$ has the eigenvalues $\mu_{i},$ $i=1,2,\ldots,m,$ we obtain ${\prod\limits_{i=1}^{m}(\Sigma_{3}-\mu_{i}I)}=\bm{0}.$ Therefore, from (5.21), we obtain $g(\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B})=\bm{0}.$ Hence, by Cayley-Hamilton theorem, the degree of the minimal polynomial of $\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B}$ is at most $m+1.$
As mentioned in [35], the degree of the minimal polynomial of a matrix and the dimension of the associated Krylov subspace are equal. Therefore, the Krylov subspace $\mathcal{K}(\mathcal{P}_{\text{RGSS-II}}^{-1}\mathcal{B},{\bm{b}})$ has dimension at most $m+1.$ Hence, the proof follows. $\blacksquare$
@@ 6 Discussion on the selection of the parameters
It is worth noting that proposed GSS, RGSS-I and RGSS-II preconditioners involve the parameters $\alpha,$ $\beta,$ $\tau$ and $\bm{\omega}.$ As a general criterion for a preconditioner to perform efficiently, it should be as close as possible to the coefficient matrix of the system [10], we find the optimal choices for the proposed GSS preconditioner by minimizing $\|\mathcal{N}_{\text{GSS}}\|_{F}=\|\mathcal{P}_{\text{GSS}}-\mathcal{B}\|_{F}.$ First, we define the function $\varphi$ by
@@
In the sequel, the GSS preconditioner is equivalently rewritten as: $\mathcal{P}_{\text{GSS}}=\bm{\omega}\widetilde{\mathcal{P}}_{\text{GSS}},$ where
$\widetilde{\mathcal{P}}_{\text{GSS}}=\left[\begin{matrix}\frac{\alpha}{\bm{\omega}}P+A&\bm{0}&B^{T}\\ \bm{0}&\frac{\beta}{\bm{\omega}}Q+D&C\\ -B&-C^{T}&\frac{\tau}{\bm{\omega}}R\end{matrix}\right]$
can be regarded as a scaled preconditioner. A preconditioner is considered efficient if it closely approximates the coefficient matrix, and for given $\alpha,\beta,\tau>0,$ $\widetilde{\mathcal{P}}_{\text{GSS}}-\mathcal{B}=\frac{1}{\bm{\omega}}\Theta\rightarrow\bm{0}$ as $\bm{\omega}\rightarrow\infty.$ Thus, the GSS preconditioner shows enhanced efficiency for large values of $\bm{\omega}.$ Similar studies apply to the RGSS-I and RGSS-II preconditioners as well. The performance of the proposed preconditioners by varying the parameters is shown in the numerical experiment section.
@@ 7 Numerical experiments
To demonstrate the effectiveness and robustness of the proposed preconditioners GSS, RGSS-I and RGSS-II over the existing ones within the Krylov subspace iterative method to solve the DSPP, in this section, we perform a few numerical experiments. We compare our proposed GSS, RGSS-I and RGSS-II PGMRES methods (abbreviated as “GSS”, “RGSS-I” and “‘RGSS-II”, respectively) with the GMRES (without preconditioning) method and PGMRES methods with block diagonal [13], block preconditioner [26], dimension splitting [8], relaxed dimension factorization [11], and shift-splitting [19] preconditioners (abbreviated as “BD”, “BP” “DS”, “RDF”, and “SS”, respectively). Moreover, we have also compared proposed methods with BD preconditioner in conjunction with minimum residual method (MIRES) (abbreviated as “BD-MINRES”). The numerical results are presented in terms of iteration counts (abbreviated as “IT”) and elapsed CPU time in seconds (abbreviated as “CPU”).
@@
or if the maximum number of iterations exceeds $5000$. Numerical experiments are conducted in MATLAB R2024a on a Windows 11 system, using an Intel(R) Core(TM) i7-10700 CPU at 2.90 GHz with 16 GB of memory.
@@
$\begin{array}[]{cc}&\underset{u,f}{\text{min}}\leavevmode\nobreak\ \frac{1}{2}\|u-\hat{u}\|^{2}_{L_{2}(\Omega)}+\frac{{\bm{\nu}}}{2}\|f\|^{2}_{L_{2}(\Omega)}\\ \text{such that }&-\nabla^{2}u=f\leavevmode\nobreak\ \text{in}\leavevmode\nobreak\ \Omega,\\ &\quad\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u=g\leavevmode\nobreak\ \text{on}\leavevmode\nobreak\ \partial\Omega,\end{array}$ (7.4)
where $u$ is the state, and $\hat{u}$ is the desired state, $f$ is the control, $\Omega=[0,1]\times[0,1]$ is the domain with the boundary $\partial\Omega,$ and $0<{{\bm{\nu}}}\ll 1$ is the regularization parameter. By discretizing (7.4) using the Galerkin finite element method and then applying Lagrange multiplier techniques, we obtain the following system:
$\left[\begin{matrix}{{\bm{\nu}}}M&\bm{0}&K^{T}\\ \bm{0}&M&-M\\ -K&M&\bm{0}\end{matrix}\right]\left[\begin{matrix}u\\ f\\ \bm{\lambda}\end{matrix}\right]=\left[\begin{matrix}\bm{0}\\ b\\ -d\end{matrix}\right],$ (7.5)
where $M\in{\mathbb{R}}^{n\times n}$ and $K\in{\mathbb{R}}^{n\times n}$ are SPD mass matrix and discrete Laplacian, respectively. Note that by setting $A={\bm{\nu}}M$, $B=K$, $C=-M$, and $\mathbf{b}=[\mathbf{0},b^{T},-d^{T}]^{T}$, (7.5) can be expressed in the form of the DSPP (1.1), where $n=m=l$.
The MATLAB code downloaded from [32], generates the linear system of the form (7.5) by using the following setup: parameters in “set_def_setup.m” are selected as def_setup.bc=‘dirichlet’, def_setup.beta=$1e-2$, def_setup.ob=1, def_setup.type=‘dist2d’ and def_setup.pow = $5$, $6$ and $7$. For these selection of def_setup.pow, size of the coefficient matrix $\mathcal{B}$ of the DSPP (1.1) is $2883,$ $11907$ and $48378,$ respectively.
Table 1: Experimental results of GMRES, BD, BD-MINRES, BP, DS, RDF, SS, GSS, IGSS, RGSS-I, IRGSS-I, RGSS-II, and IRGSS-II PGMRES methods for Example 1 when ${{\bm{\nu}}}=0.1$
@@
Table 2: Experimental results of GMRES, BD, BD-MINRES, BP, DS, RDF, SS, GSS, IGSS, RGSS-I, IRGSS-I, RGSS-II, and IRGSS-II PGMRES methods for Example 1 when ${{\bm{\nu}}}=0.001$